OpenStudy (anonymous):

I have a very stupid question. I'm quite old and haven't done algebra since more than 15 years... now I stumbled upon this problem which I cannot solve. I feel ashamed. Can someone help me ? Here it is: y = x + 2 z = x +3 y = 2z + 1 Note: I don't want the answer, only a clear explanation of steps to use to resolve the problem. Thank you very much. :)

OpenStudy (anonymous):

I'm 32, haha...

OpenStudy (anonymous):

Old related to how long it's been since I last checked an algebra problem that is. ;)

OpenStudy (radar):

You could use the elimination method.

OpenStudy (anonymous):

I don't have a clue of whatever method whatsoever. I want a clear explanation with steps. :) No answer please.

OpenStudy (unklerhaukus):

y = 2z + 1 y = 2(x +3) + 1 y = 2((y-2) +3) + 1 y=2(y+1)+1 y=2y+3 y=-3/2

OpenStudy (anonymous):

Oh my, I wasn't expecting a solution so soon, haha... :) I'll study those answers carefully. Thank you very much to all contributors, I love you, haha! :)

OpenStudy (radar):

You have a system of three equations and you have three unknowns so theoretically this system can be solved. It is worth trying lol. Take these two y=x+2 and y=2z+1 since y=y then we can say x+2=2z+1 obtaining a new equation x=2z-1.

OpenStudy (anonymous):

Thank you very much everyone.

OpenStudy (turingtest):

UnkleRhaukus used substitution y = x + 2 (1) z = x +3 (2) y = 2z + 1 (3) I would sub that expression for z in equation (2) into equation (3), which will give y in terms of x. between that and equation (1) you should be able to get a number value for x or y. From there on it is a system of two variables, so you can probably take it from there. Good luck!

OpenStudy (radar):

Now I want to come up with another equation that has only the variables x and z how about z=x+3 (given) from that we get x=z-3. Now we try to eliminate a variable. x=z-3 and x=2z-1 x=2z-1 subtract 0=-z-4 z=-4 from there you substitute finding the others. You manipulate the equations to eliminate a variable, then solve for an unknown and the follow through with substitution. By the way I am 73, and that it "senior"

OpenStudy (radar):

*it is